# A Fully Data-Driven Value Iteration for Stochastic LQR: Convergence, Robustness and Stability

### v2

@@ 6. An Application to Dynamic Portfolio Allocation
In this section, we apply the R-LSVI algorithm to a practical financial problem, referred to as dynamic portfolio allocation, using a modeling framework similar to Gârleanu and Pedersen [21]. This model admits a standard stochastic LQR representation [3, 46], with the state vector $x_{t}\in\mathbb{R}^{2N+M}$ that is the concatenation of the portfolio weight in each asset at time t1t-1 ($w_{t-1}\in\mathbb{R}^{N}$), mean-reverting strategy signals ($f_{t}\in\mathbb{R}^{M}$) and assets returns between t1t-1 ($r_{t}\in\mathbb{R}^{N}$) and the control input $u_{t}\in\mathbb{R}^{N}$ representing the change in portfolio weights at time $t$ ($u_{t}\triangleq w_{t}-w_{t-1}$). In this model we assume the asset returns, $r_{t+1}$, are a linear combination of the mean-reverting signal $f_{t}$ leading to the linear dynamic
@@
where $\Pi\in\mathbb{R}^{N\times M}$ is the matrix of factor loadings, $\Phi\in\mathbb{R}^{M\times M}$ is the matrix of the mean-reversion coefficients with $\rho(I_{M}-\Phi)<1$, $\Sigma\in\mathbb{S}^{N}_{++}$, $\Omega\in\mathbb{S}^{M}_{++}$, and $\{\epsilon^{r}_{t}\}$ and $\{\epsilon^{f}_{t}\}$ are the covariance matrix of the white processes representing return noise and factors innovations.
@@
Regarding cost misspecification, the left panel of Figure 4 shows that the nominal model-based control slightly outperforms the non-model-based methods for $\alpha\leq 2$, in contrast to the previous example, where this occurred only at $\alpha=2$. In the right panel, however, when $\alpha>2$, the nominal approach fails to produce a stable control gain, leading to a dramatic increase in final cost. In contrast, the non-model-based methods maintain robust performance, achieving $100\%$ stability with final costs of comparable magnitude.
Figure 4. Convergence and stability of the nominal VI, O-LSPI and R-LSVI algorithms on the dynamic portfolio allocation problem with non-quadratic cost. (Left) Costs as the exponent $\alpha$ varies. Dashed lines represent the median relative error, with the shaded region covering the $25^{th}$ to $75^{th}$ percentiles, estimated from 100100 trajectories. (Right) Frequency of stabilizing controllers found by the algorithms. Only costs corresponding to stable control gains are plotted in the left panel.
@@ Appendix A Auxiliary Results
$K_{i}^{j}$ $\triangleq(R+B^{\top}P_{i}^{j}B)^{-1}B^{\top}P_{i}^{j}A$ (92)
$A_{i}^{j}$ $\triangleq A-BK_{i}^{j}$ (93)
$R_{i}^{j}$ $\triangleq R+B^{\top}P_{i}^{j}B$ (94)
@@
    $\mathcal{Q}\left(P\right)\succeq\mathcal{Q}\left(0_{n\times n}\right)\succeq 0$.
@@
    If $\lVert\Delta\mathcal{Q}\left(P\right)\rVert_{2}<d$ then $\hat{\mathcal{Q}}\left(P\right)\in\mathbb{S}^{n+m}_{+}$.
@@
(i) We have that $\mathcal{Q}\left(0_{n\times n}\right)=\begin{bmatrix}S&0\\ 0&R\end{bmatrix}\succeq 0$, as $S\in\mathbb{S}^{n}_{+}$ and $R\in\mathbb{S}^{m}_{++}$. Hence,
@@
$\lambda_{\mathrm{min}}\left(\hat{\mathcal{Q}}(P)\right)$ $\geq\lambda_{\mathrm{min}}\left(\mathcal{Q}\left(P\right)\right)-\lambda_{\mathrm{max}}\left(\Delta\mathcal{Q}\left(P\right)\right)$ (104)
@@
$\geq 0.$ (106)
(iii) Recall that $\hat{\mathcal{Q}}\left(P\right)]_{uu}=R+B^{\top}PB$, and hence
$\lambda_{\mathrm{min}}\left([\hat{\mathcal{Q}}\left(P\right)]_{uu}\right)$ $\geq\lambda_{\mathrm{min}}\left(R\right)-\lambda_{\mathrm{max}}\left([\Delta\mathcal{Q}\left(P\right)]_{uu}\right)$ (107)
$\geq\lambda_{\mathrm{min}}\left(R\right)-\lambda_{\mathrm{max}}\left(\Delta\mathcal{Q}\left(P\right)\right)$ (108)
@@
as $\lambda_{\mathrm{min}}\left(R\right)>0$.
@@
$\Delta{P}_{i+1}$ $=(A_{i}^{2})^{\top}[\Delta{P}_{i}-\Delta{P}_{i}B(R+B^{\top}{P}_{i}B)^{-1}B^{\top}\Delta{P}_{i}]A_{i}^{2}+\Delta S_{i}$ (120)
@@
$\Delta{P}_{i+1}$ $=(A_{i}^{2})^{\top}\Delta{P}_{i}A_{i}^{2}+\Delta S_{i}$ (125)
$\quad-(K_{i}^{1}-K_{i}^{2})^{\top}(R+B^{\top}P_{i}^{1}B)(K_{i}^{1}-K_{i}^{2}).$
@@
$\Delta{P}_{i+1}$ $=(A_{i}^{2})^{\top}\Delta{P}_{i}A_{i}^{2}+\Delta S_{i}$ (126)
$\quad-(A_{i}^{2})^{\top}\Delta{P}_{i}B(R+B^{\top}{P}_{i}B)^{-1}B^{\top}\Delta{P}_{i}A_{i}^{2}.$
@@ Appendix B Proof of Proposition 1
with $L=A-BK_{c}$ and $D=(B\Sigma_{\eta}^{1/2}\penalty 10000\ C)$. Without loss of generality, we assume that $(L,D)$ is controllable. Since $x_{0}=0_{n}$, each uncontrollable state component is identically equal to zero for all $t$. Its invariant distribution is therefore the point mass $\delta_{0}$.
@@ Appendix C Exact VI
The proof is by induction. Note that $\lVert\Delta\rVert_{\infty}=\sup\{\lVert\Delta_{i}\rVert:i\in\mathbb{Z}_{+}\}\leq\lambda_{\mathrm{min}}\left(S\right)$ implies that $S+\Delta_{i}\succ 0$ for all $i\in\mathbb{Z}_{+}$. Suppose $\hat{P}_{i}\succeq 0$ for some $i\in\mathbb{Z}_{+}$. Since
$\hat{P}_{i+1}$ $=\mathcal{R}(\hat{P}_{i},S+\Delta_{i})$ (134)
$=A^{\top}\left(\hat{P}_{i}-\hat{P}_{i}B(R+B^{\top}\hat{P}_{i}B)^{-1}B^{\top}\hat{P}_{i}\right)A+S+\Delta_{i}$
@@
If $\Delta P_{0}\triangleq P_{0}-P^{*}\in\mathcal{P}\triangleq\{P\in\mathbb{S}^{n}:\lVert P\rVert\leq\frac{1-\lVert A^{*}\rVert}{\lVert BR^{-1}B^{\top}\rVert\lVert A^{*}\rVert}\}$, then $\lVert\Delta P_{i}\rVert\leq\lVert A^{*}\rVert^{i}\lVert\Delta P_{0}\rVert$ for all $i\in\mathbb{Z}_{+}$.
@@
$\lVert\Delta{P}_{i+1}\rVert$ $\leq\lVert A^{*}\rVert\lVert\Delta{P}_{i}\rVert(1+\lVert BR^{-1}B^{\top}\rVert\lVert\Delta{P}_{i}\rVert)\lVert A^{*}\rVert.$ (138)
If $\Delta P_{i}\in\mathcal{P}$, we have $\lVert\Delta{P}_{i+1}\rVert\leq\lVert A^{*}\rVert\lVert\Delta{P}_{i}\rVert$, and the proof is complete. ∎
@@ C.1. Proof of Proposition 2
$\Delta{P}_{i+1}$ $=(A_{i}^{2})^{\top}\Delta{P}_{i}A_{i}^{2}+\Delta S_{i}$ (139)
$\quad-(A_{i}^{2})^{\top}\Delta{P}_{i}B(R+B^{\top}P_{i}^{1}B)^{-1}B^{\top}\Delta{P}_{i}A_{i}^{2}$
$=(A_{i}^{2})^{\top}\Delta{P}_{i}A_{i}^{2}+\Delta S_{i}$ (140)
$\quad-(A_{i}^{2})^{\top}\Delta{P}_{i}B(R^{2}_{i}+B^{\top}\Delta{P}_{i}B)^{-1}B^{\top}\Delta{P}_{i}A_{i}^{2}$
$=\Delta S_{i}+(A_{i}^{2})^{\top}[\Delta{P}_{i}$ (141)
$\quad-\Delta{P}_{i}B(R^{2}_{i}+B^{\top}\Delta{P}_{i}B)^{-1}B^{\top}\Delta{P}_{i}]A_{i}^{2}.$
By applying Lemma 4, we have
$\Delta{P}_{i}-\Delta{P}_{i}B(R^{2}_{i}+B^{\top}\Delta{P}_{i}B)^{-1}B^{\top}\Delta{P}_{i}\succeq 0\,.$ (142)
Therefore, we conclude that $\Delta{P}_{i+1}\succeq 0$, i.e. $P_{i+1}^{1}\succeq P_{i+1}^{2}$.
@@ C.2. Proof of Proposition 3
where $K(P)=(R+B^{T}PB)^{-1}B^{\top}PA$. Therefore, for $P=P^{*}(\delta)$ with $\delta\in(-\lambda_{\mathrm{min}}\left(S\right),\infty)$, $A\left(P^{*}(\delta)\right)$ is Hurwitz (see, [44, Section 2.4] for details). Since $\lVert A\left(P^{*}(\delta)\right)\rVert<1$, the left hand side of (147) is invertible. Consequently, by the implicit function theorem $P^{*}(\delta)$ is continuously differentiable with respect to $\delta$.
@@ C.3. Proof of Theorem 2
Let $\Delta P_{i}=P_{i}-P^{*}$. Then, by Lemma 6, we have
$\Delta{P}_{i+1}$ $=(A^{*})^{\top}\Delta{P}_{i}A^{*}-(A^{*})^{\top}\Delta{P}_{i}B(R^{*}+B^{\top}\Delta{P}_{i}B)^{-1}B^{\top}\Delta{P}_{i}A^{*}\,.$ (149)
@@
$\Delta{P}_{0}\succeq 0$: First, we show by induction that if $\Delta{P}_{0}\succeq 0$ then $\Delta{P}_{i}\succeq 0$ for all $i\in\mathbb{Z}_{+}$. Indeed, assume $\Delta{P}_{i}\succeq 0$ for some $i\in\mathbb{Z}_{+}$. Then by Lemma 6
Since $\Delta{P}_{i}\succeq[\Delta{P}_{i}-\Delta{P}_{i}B(R^{*}+B^{\top}\Delta{P}_{i}B)^{-1}B^{\top}\Delta{P}_{i}]\succeq 0$ by Lemma 8, we have that $\Delta{P}_{i+1}\succeq 0$. Taking the norm of both sides of (150) results in
@@
$\Delta{P}_{0}$ is indefinite: Obviously, $\lambda_{\mathrm{min}}\left(\Delta{P}_{0}\right)I_{n}\preceq\Delta{P}_{0}\preceq\lambda_{\mathrm{max}}\left(\Delta{P}_{0}\right)I_{n}$. Let $\{P^{+}_{i}\}_{i\in\mathbb{Z}_{+}}$ and $\{P^{-}_{i}\}_{i\in\mathbb{Z}_{+}}$ be the sequences generated by the exact VI with $P^{+}_{0}=\lambda_{\mathrm{max}}\left(\Delta{P}_{0}\right)I_{n}+P^{*}$ and $P^{-}_{0}=\lambda_{\mathrm{min}}\left(\Delta{P}_{0}\right)I_{n}+P^{*}$, respectively. Since $P_{0}^{+}-P^{*}\succeq 0$, it follows from Case ① that $\lVert P_{i}^{+}-P^{*}\rVert_{2}\leq\beta(\lambda_{\mathrm{max}}\left(P_{0}-P^{*}\right),i)$. In addition, since $P^{-}_{0}-P^{*}\preceq 0$, by Case ②, we have that $\lVert P_{i}^{-}-P^{*}\rVert_{2}\leq\beta(-\lambda_{\mathrm{min}}\left(P_{0}-P^{*}\right),i)$. Since $P_{0}^{-}\preceq P_{0}\preceq P_{0}^{+}$, it follows from Proposition 2, $P_{i}^{-}-P^{*}\preceq P_{i}-P^{*}\preceq P_{i}^{+}-P^{*}$ for all $i\in\mathbb{Z}_{+}$. Consequently, $\lambda_{\mathrm{min}}\left(P_{i}^{-}-P^{*}\right)\leq\lambda_{\mathrm{min}}\left(P_{i}-P^{*}\right)\leq\lambda_{\mathrm{max}}\left(P_{i}-P^{*}\right)\leq\lambda_{\mathrm{max}}\left(P_{i}^{+}-P^{*}\right)$. Hence, we have
@@
$\leq\max\{\beta(\lambda_{\mathrm{max}}\left(P_{0}-P^{*}\right),i),\beta(-\lambda_{\mathrm{min}}\left(P_{0}-P^{*}\right),i)\}$

### v3

@@ 6. An Application to Dynamic Portfolio Allocation
In this section, we apply the R-LSVI algorithm to a practical financial problem, referred to as dynamic portfolio allocation, using a modeling framework similar to Gârleanu and Pedersen [21]. This model admits a standard stochastic LQR representation [3, 46], with the state vector $x_{t}\in\mathbb{R}^{2N+M}$ that is the concatenation of the portfolio weights in each asset at time t1t-1 ($w_{t-1}\in\mathbb{R}^{N}$), mean-reverting strategy signals ($f_{t}\in\mathbb{R}^{M}$) and asset returns between [t1,t)[t-1,t) ($r_{t}\in\mathbb{R}^{N}$) and the control input $u_{t}\in\mathbb{R}^{N}$ representing the change in portfolio weights at time $t$ ($u_{t}\triangleq w_{t}-w_{t-1}$). In this model we assume the asset returns, $r_{t+1}$, are a linear combination of the mean-reverting signal $f_{t}$ leading to the linear dynamic
@@
where $\Pi\in\mathbb{R}^{N\times M}$ is the matrix of factor loadings, $\Phi\in\mathbb{R}^{M\times M}$ is the matrix of the mean-reversion coefficients with $\rho(I_{M}-\Phi)<1$, $\Sigma\in\mathbb{S}^{N}_{++}$, $\Omega\in\mathbb{S}^{M}_{++}$, and $\{\epsilon^{r}_{t}\}$ and $\{\epsilon^{f}_{t}\}$ are the covariance matrix of the white noise processes representing return noise and factors innovations.
@@
Regarding cost misspecification, the left panel of Figure 4 shows that the nominal model-based control slightly outperforms the non-model-based methods for ${\kappa}\leq 2$, in contrast to the previous example, where this occurred only at ${\kappa}=2$. In the right panel, however, when ${\kappa}>2$, the nominal approach fails to produce a stable control gain, leading to a dramatic increase in final cost. In contrast, the non-model-based methods maintain robust performance, achieving $100\%$ stability with final costs of comparable magnitude.
Figure 4. Convergence and stability of the nominal VI, O-LSPI and R-LSVI algorithms on the dynamic portfolio allocation problem with non-quadratic cost. (Left) Costs as the exponent $\kappa$} varies. Dashed lines represent the median relative error, with the shaded region covering the $25^{th}$ to $75^{th}$ percentiles, estimated from 100100 trajectories. (Right) Frequency of stabilizing controllers found by the algorithms. Only costs corresponding to stable control gains are plotted in the left panel.
@@ Appendix A Auxiliary Results
$K_{i}^{j}\triangleq(R+B^{\top}P_{i}^{j}B)^{-1}B^{\top}P_{i}^{j}A,\quad A_{i}^{j}\triangleq A-BK_{i}^{j},\quad R_{i}^{j}\triangleq R+B^{\top}P_{i}^{j}B$ (93)
@@
    $\mathcal{Q}\left(P\right)\succeq\mathcal{Q}\left(0_{n\times n}\right)\succ 0$.
@@
    If $\lVert\Delta\mathcal{Q}\left(P\right)\rVert_{2}<d$ then $\hat{\mathcal{Q}}\left(P\right)\in\mathbb{S}^{n+m}_{++}$.
@@
(i) We have that $\mathcal{Q}\left(0_{n\times n}\right)=\begin{bmatrix}S&0\\ 0&R\end{bmatrix}\succ 0$, as $S\in\mathbb{S}^{n}_{++}$ and $R\in\mathbb{S}^{m}_{++}$. Hence,
@@
$\lambda_{\mathrm{min}}\left(\hat{\mathcal{Q}}(P)\right)$ $\geq\lambda_{\mathrm{min}}\left(\mathcal{Q}\left(P\right)\right)-\lVert\Delta\mathcal{Q}\left(P\right)\rVert_{2}$ (103)
@@
$>0.$ (105)
(iii) Recall that $\hat{\mathcal{Q}}\left(P\right)]_{uu}=R+B^{\top}PB{+[\Delta\mathcal{Q}\left(P\right)]_{uu}}$, hence
$\lambda_{\mathrm{min}}\left([\hat{\mathcal{Q}}\left(P\right)]_{uu}\right)$ $\geq\lambda_{\mathrm{min}}\left(R\right)-{\lVert[\Delta\mathcal{Q}\left(P\right)]_{uu}\rVert_{2}}$ (106)
$\geq\lambda_{\mathrm{min}}\left(R\right)-{\lVert\Delta\mathcal{Q}\left(P\right)\rVert_{2}}$ (107)
@@
as $\lambda_{\mathrm{min}}\left(R\right)>0$. ∎
@@
$\Delta{P}_{i+1}$ $=(A_{i}^{2})^{\top}[\Delta{P}_{i}-\Delta{P}_{i}B(R+B^{\top}{P}^{1}_{i}B)^{-1}B^{\top}\Delta{P}_{i}]A_{i}^{2}+\Delta S_{i}$ (119)
@@
$\Delta{P}_{i+1}$ $=(A_{i}^{2})^{\top}\Delta{P}_{i}A_{i}^{2}+\Delta S_{i}-(K_{i}^{1}-K_{i}^{2})^{\top}(R+B^{\top}P_{i}^{1}B)(K_{i}^{1}-K_{i}^{2}).$ (124)
@@
$\Delta{P}_{i+1}$ $=(A_{i}^{2})^{\top}\Delta{P}_{i}A_{i}^{2}+\Delta S_{i}-(A_{i}^{2})^{\top}\Delta{P}_{i}B(R+B^{\top}{P}^{1}_{i}B)^{-1}B^{\top}\Delta{P}_{i}A_{i}^{2}.$ (125)
@@ Appendix B Proof of Proposition 1
with $L=A-BK_{c}$ and $D=(B\Sigma_{\eta}^{1/2}\penalty 10000\ C)$. Without loss of generality, we assume that $(L,D)$ is controllable. Since $x_{0}=0_{n}$, each uncontrollable state component is identically equal to zero for all $t$. Their invariant distribution is therefore the point mass $\delta_{0}$.
@@ Appendix C Exact VI
The proof is by induction. Note that $\lVert\Delta\rVert_{\infty}=\sup\{\lVert\Delta_{i}\rVert,i\in\mathbb{Z}_{+}\}<\lambda_{\mathrm{min}}\left(S\right)$ implies that $S+\Delta_{i}\succ 0$ for all $i\in\mathbb{Z}_{+}$. Suppose $\hat{P}_{i}\succeq 0$ for some $i\in\mathbb{Z}_{+}$. Since
$\hat{P}_{i+1}$ $=\mathcal{R}(\hat{P}_{i},S+\Delta_{i})=A^{\top}\left(\hat{P}_{i}-\hat{P}_{i}B(R+B^{\top}\hat{P}_{i}B)^{-1}B^{\top}\hat{P}_{i}\right)A+S+\Delta_{i}$ (133)
@@
If $\Delta P_{0}\triangleq P_{0}-P^{*}\in\mathcal{P}\triangleq\{P\in\mathbb{S}^{n}:\lVert P\rVert_{{2}}\leq\frac{1-\lVert A^{*}\rVert_{{2}}}{\lVert BR^{-1}B^{\top}\rVert_{{2}}\lVert A^{*}\rVert_{{2}}}\}$, then for all $i\in\mathbb{Z}_{+}$
$\lVert\Delta P_{i}\rVert_{{2}}\leq\lVert A^{*}\rVert_{{2}}^{i}\lVert\Delta P_{0}\rVert_{{2}}\,.$ (135)
@@
$\lVert\Delta{P}_{i+1}\rVert_{{2}}$ $\leq\lVert A^{*}\rVert_{{2}}\lVert\Delta{P}_{i}\rVert_{{2}}(1+\lVert BR^{-1}B^{\top}\rVert_{{2}}\lVert\Delta{P}_{i}\rVert_{2})\lVert A^{*}\rVert_{{2}}.$ (138)
If $\Delta P_{i}\in\mathcal{P}$, we have $\lVert\Delta{P}_{i+1}\rVert_{{2}}\leq\lVert A^{*}\rVert_{{2}}\lVert\Delta{P}_{i}\rVert_{{2}}$, and the proof is complete. ∎
@@ C.1. Proof of Proposition 2
$\Delta{P}_{i+1}$ $=\Delta S_{i}+(A_{i}^{2})^{\top}[\Delta{P}_{i}-\Delta{P}_{i}B(R^{2}_{i}+B^{\top}\Delta{P}_{i}B)^{-1}B^{\top}\Delta{P}_{i}]A_{i}^{2}.$ (139)
By applying Lemma 4, we have that $\Delta{P}_{i+1}\succeq 0$, i.e. $P_{i+1}^{1}\succeq P_{i+1}^{2}$.
@@ C.2. Proof of Proposition 3
where $K(P)=(R+B^{T}PB)^{-1}B^{\top}PA$. Therefore, for $P=P^{*}(\delta)$ with $\delta\in(-\lambda_{\mathrm{min}}\left(S\right),\infty)$, $A\left(P^{*}(\delta)\right)$ is Schur stable (see, [44, Section 2.4] for details). Since $\lVert A\left(P^{*}(\delta)\right)\rVert<1$, the left hand side of (144) is invertible. Consequently, by the implicit function theorem $P^{*}(\delta)$ is continuously differentiable with respect to $\delta$.
@@ C.3. Proof of Theorem 2
Let $\Delta P_{i}=P_{i}-P^{*}$. Then, Using Lemma 6 and that $R+B^{\top}P_{i}B=R^{*}+B^{\top}\Delta{P}_{i}B$, we obtain
@@
$\Delta{P}_{0}\succeq 0$: First, we show by induction that if $\Delta{P}_{0}\succeq 0$ then $\Delta{P}_{i}\succeq 0$ for all $i\in\mathbb{Z}_{+}$. Indeed, assume $\Delta{P}_{i}\succeq 0$ for some $i\in\mathbb{Z}_{+}$. Since $\Delta{P}_{i}\succeq[\Delta{P}_{i}-\Delta{P}_{i}B(R^{*}+B^{\top}\Delta{P}_{i}B)^{-1}B^{\top}\Delta{P}_{i}]\succeq 0$ by Lemma 4, we have that $\Delta{P}_{i+1}\succeq 0$. Taking the norm of both sides of (146) results in
@@
$\Delta{P}_{0}$ is indefinite: Obviously, $\lambda_{\mathrm{min}}\left(\Delta{P}_{0}\right)I_{n}\preceq\Delta{P}_{0}\preceq\lambda_{\mathrm{max}}\left(\Delta{P}_{0}\right)I_{n}$. Let $\{P^{+}_{i}\}_{i\in\mathbb{Z}_{+}}$ and $\{P^{-}_{i}\}_{i\in\mathbb{Z}_{+}}$ be the sequences generated by the exact VI with $P^{+}_{0}=\lambda_{\mathrm{max}}\left(\Delta{P}_{0}\right)I_{n}+P^{*}$ and $P^{-}_{0}=\lambda_{\mathrm{min}}\left(\Delta{P}_{0}\right)I_{n}+P^{*}$, respectively. Since $P_{0}^{+}-P^{*}\succeq 0$, it follows from Case ① that $\lVert P_{i}^{+}-P^{*}\rVert_{2}\leq\beta(\lambda_{\mathrm{max}}\left(P_{0}-P^{*}\right),i)$. In addition, since $P^{-}_{0}-P^{*}\preceq 0$, by Case ②, we have that $\lVert P_{i}^{-}-P^{*}\rVert_{2}\leq$ $\beta(-\lambda_{\mathrm{min}}\left(P_{0}-P^{*}\right),i)$. Since $P_{0}^{-}\preceq P_{0}\preceq P_{0}^{+}$, it follows from Proposition 2,
$P_{i}^{-}-P^{*}\preceq P_{i}-P^{*}\preceq P_{i}^{+}-P^{*}$ (152)
for all $i\in\mathbb{Z}_{+}$. Consequently, $\lambda_{\mathrm{min}}\left(P_{i}^{-}-P^{*}\right)\leq\lambda_{\mathrm{min}}\left(P_{i}-P^{*}\right)\leq 0\leq\lambda_{\mathrm{max}}\left(P_{i}-P^{*}\right)\leq\lambda_{\mathrm{max}}\left(P_{i}^{+}-P^{*}\right)$. Hence, we have
@@
$\leq\max\{\beta(-\lambda_{\mathrm{min}}\left(P_{0}-P^{*}\right),i),\beta(\lambda_{\mathrm{max}}\left(P_{0}-P^{*}\right),i)\}$